### v1

Abstract
Given a sequence of tokens, such as words, the task of next-token prediction is to predict the next-token conditional probability distribution. Decoder-only transformers have become effective models for this task, but their properties are still not fully understood. In particular, the largest number of distinct context sequences that a decoder-only transformer can interpolate next-token distributions for has not been established. To fill this gap, we prove upper and lower bounds on this number, which are equal up to a multiplicative constant. We prove these bounds in the general setting where next-token distributions can be arbitrary as well as the empirical setting where they are calculated from a finite number of document sequences. Our lower bounds are for one-layer transformers and our proofs highlight an important injectivity property satisfied by self-attention. Furthermore, we provide numerical evidence that the minimal number of parameters for memorization is sufficient for being able to train the model to the entropy lower bound.
1 Introduction
Large language models have become unquestionably effective at certain language tasks, but are they really learning? And if they are, are they learning in ways that we expect; are they learning in a way that can even be tractably unraveled? A researcher with some intuitions about how language works may invent an effective language model, but that does not mean the intuitions are correct. In this paper, we will prove, without recourse to any intuitions, that the now ubiquitous decoder-only transformer model is capable of interpolating arbitrary data sets, in the next-token prediction setting, using the minimal number of parameters required. In fact, it can even do this with an extremely simplified architecture which bares more resemblance to polynomial interpolation than to natural language processing. To get there, we will first give a brief overview of relevant concepts.
Ten years after Bengio et al (2000) introduced neural probabilistic language models, Tomáš Mikolov constructed a language model that takes a sequence of tokens (e.g. words) as input, embeds the tokens into a vector space using a Skip-gram model Mikolov et al (2013), then inputs the sequence of token vectors to a recurrent neural network (RNN) (Mikolov, 2012). Three years later, the attention mechanism was developed for combining a sequence of vectors into a single vector (Bahdanau et al, 2015). One year later, self-attention was developed as a modification to attention (Cheng et al, 2016). One year after that, Vaswani et al (2017) recognized that self-attention made it unnecessary to use a RNN and so replaced the RNN with a feedforward neural network (FNN), introducing the modern transformer model. This gives a brief timeline of the progress leading to modern transformers. But language models have been around for much longer.
75 years ago, Claude Shannon introduced the concept of the entropy of language and used the predictability of language to study it (Shannon, 1948, 1951). The entropy of a discrete probability distribution, $P$, over $\omega$ possible events, is the quantity $-\sum_{\gamma=1}^{\omega}P(\gamma)\log(P(\gamma))$, and it satisfies certain properties such that it intuitively captures “how much ‘choice’ is involved in the selection of the event or of how uncertain we are of the outcome” (Shannon, 1948). If we are trying to fit a probability distribution, $Q_{\theta}$, parameterized by the vector $\theta\in\mathbb{R}^{k}$, to $P$, then we may use the cross-entropy loss, $-\sum_{\gamma=1}^{\omega}P(\gamma)\log(Q_{\theta}(\gamma))$, which is greater than or equal to the entropy of $P$ by Gibb’s inequality. More generally, the entropy of a probability distribution, $P$, over sequences of $\omega$ possible events, is $-\sum_{t=1}^{\infty}\sum_{\alpha\in[\omega]^{t}}\sum_{\gamma=1}^{\omega}P(\alpha,\gamma)\log(P(\alpha,\gamma\mid\alpha))$, and it lower bounds the cross-entropy $-\sum_{t=1}^{\infty}\sum_{\alpha\in[\omega]^{t}}\sum_{\gamma=1}^{\omega}P(\alpha,\gamma)\log(Q_{\theta}(\alpha,\gamma\mid\alpha))$. The idea of “next-token prediction” is to tune $\theta\in\mathbb{R}^{k}$ to minimize the cross-entropy.
Of course, in practice, $P$ is replaced with the empirical next-token distribution $\hat{P}$. For example, given a data set of $N$ document sequences $\beta^{1},\ldots,\beta^{N}\in[\omega]^{T+1}$, the cross-entropy loss is proportional to
$-\sum_{i=1}^{N}\sum_{t=1}^{T}\log(Q_{\theta}(\beta_{1}^{i},\ldots,\beta_{t+1}^{i}\mid\beta_{1}^{i},\ldots,\beta_{t}^{i}))$. If $\hat{P}(\alpha)$ is nonzero for $n$ $\alpha$’s in $\cup_{t=1}^{T}[\omega]^{t}$, i.e. if the document sequences have $n$ unique contexts, $\alpha^{i}=(\beta_{1}^{j},\ldots,\beta_{t}^{j})$, then we show, in Theorem 1, that $\approx n\omega$ parameters is necessary for the infimum of the cross-entropy loss to equal the entropy of $\hat{P}$. Then, in Theorem 5, we show that when $Q_{\theta}$ is a one-layer decoder-only transformer, $\approx n\omega$ parameters is actually sufficient. Interestingly, this requires that the number of feedforward neural-network (FNN) neurons be greater than or equal to $n$, which means that the FNN is able to smoothly interpolate the output vectors from the self-attention sub-layer (Bubeck and Sellke, 2021).
A one-layer decoder-only transformer model has three sub-layers: the embedding, the (multi-head) self-attention, and the FNN. We will show that this model has optimal interpolation power by first showing that it already does with embedding dimension 1 and only one self-attention head. In this setting, the token embeddings and positional embeddings are scalars. Given an input sequence, the self-attention sub-layer outputs a scalar as well. In Theorem 4, we show that it does so injectively, which allows the model to exploit the interpolation power of the FNN. This leads us to consider simple alternatives to self-attention that satisfy the same injectivity property. As it turns out, taking the weighted average of embedded token sequences, which we call “token averaging,” works as well, as we show in Theorem 6. Thus, replacing self-attention with token averaging in the one-layer decoder-only transformer model, results in another model with optimal memory capacity: Theorem 7.
The memory capacity, the largest $n$ such that a model can interpolate $n$ generic input-output pairs, is a classic topic in learning theory (Cover, 1965), but it is only recently, with the shift in perspective from the classical bias-variance trade-off to the double descent curve of Belkin et al (2019), that going beyond the interpolation threshold has been viewed as a benefit to generalization. Belkin et al (2019) argue that when there is an abundance of interpolating solutions, simple algorithms such as gradient descent tend to choose the smoothest interpolating solution, which is a form of Occam’s razor and therefore may be at the heart of the ability for highly expressive models to generalize beyond the data sets they are trained on. Thus, it is even reasonable to construct a model by thinking about the simplest way to achieve optimal memory capacity, so it is especially important to investigate the memory capacity of models which are already used in practice.
1.1 Results
We prove upper and lower bounds on the memory capacity of one-layer decoder-only transformers for next-token prediction in two settings: the general setting where the next-token distribution conditioned on a context sequence can be arbitrary, and the empirical setting where it is calculated from a finite number of document sequences. Let $k$ be the number of model parameters and $n$ be the number of distinct context sequences. Then our upper and lower bounds are summarized in Table 1. The particular constant hidden by $\Omega(\cdot)$ is given in Equation (3).
Table 1:
Memory capacity bounds for transformers
General setting Empirical setting
Upper bounds $\frac{k}{\omega-1}$, Lemma 1 $\left(2+\frac{1}{\omega-1}\right)\frac{k}{\omega-1}+1$, Theorem 1
Lower bounds $\Omega\left(\frac{k}{\omega-1}\right)$, Theorem 5 $\Omega\left(\frac{k}{\omega-1}\right)$, Lemma 2 applied to Theorem 5
We prove the memory capacity lower bound in Theorem 5 by reducing to the one-head setting with embedding dimension 1. In this setting, the self-attention sub-layer outputs scalars to the FNN sub-layer. Given $n$ distinct context sequences, let $x\in\mathbb{R}^{n}$ denote the $n$ scalar outputs. Then the model has the form $V^{\top}\psi(wx^{\top})$ where $\psi$ is the activation function of the FNN. If $\psi(wx^{\top})$ has rank $n$, then we can solve for $V$ given any target $Y\in\mathbb{R}^{\omega\times n}$. Thus, the memory capacity of $V^{\top}\psi(wx^{\top})$ comes down to the rank of $\psi(wx^{\top})$.
In Theorem 2, we show that if $\psi$ is a polynomial with $k$ monomials and $b\in\mathbb{R}^{n}$ has entries which are nonzero and distinct, then the rank and Kruskal rank of $\psi(ab^{\top})$ are $\min\{m,n,k\}$ for generic $a\in\mathbb{R}^{m}$. In Theorem 3, we show that if $\psi$ is real analytic at zero and not a polynomial there, and $b\in\mathbb{R}^{n}$ has entries which are nonzero and distinct, then the rank and Kruskal rank of $\psi(ab^{\top})$ are $\min\{m,n\}$ for generic $a$ in a neighborhood of zero in $\mathbb{R}^{m}$.
Thus, the memory capacity of $V^{\top}\psi(wx^{\top})$ comes down to whether or not the entries of $x$ are nonzero and distinct. To show that they are, we show, in Theorem 4, that self-attention maps injectively into \$\0\mathbb{R}\backslash\{0\}$. This also leads us to consider token averaging as a simple alternative to self-attention. In Theorem 6, we show that token averaging also maps injectively into \$\0\mathbb{R}\backslash\{0\}$, which leads to an optimal memory capacity lower bound, Theorem 7, for the composition of token averaging and a FNN.
1.2 Related work
The memory capacity of a machine learning model with $k$ parameters is the largest $n$ such that it can interpolate $n$ generic input-output pairs (Cover, 1965). Baum (1988) showed that a two-layer neural network with Heaviside activation has memory capacity at least $\approx k$ where outputs are in {±$\{\pm 1\}$. Yun et al (2019) proved a similar result but with rectified linear unit (ReLU) activation. Then, a year later, Bubeck et al (2020) showed that this actually applies more generally for outputs in $\mathbb{R}$. Here, because $\mathbb{R}^{k}$ and $\mathbb{R}^{n}$ are both manifolds, $k\geq n$ is necessary to interpolate $n$ generic input-output pairs, and so the memory capacity result is optimal. Four years later, Madden and Thrampoulidis (2024) proved a similar result but for arbitrary activation functions that are real analytic at a point and, if reducing to a polynomial there, have sufficiently high degree. Thus, the optimal memory capacity of two-layer neural networks has been well established.
In the setting of next-token prediction, inputs are context sequences in $\cup_{t=1}^{\infty}[\omega]^{t}$ and outputs are in $\Delta^{\omega-1}$, the unit simplex in $\mathbb{R}^{\omega}$. An output is viewed as a prediction of the distribution of the next-token conditioned on the context sequence. While Kim et al (2023), Mahdavi et al (2024), and Kajitsuka and Sato (2024) prove memory capacity results about transformers, only Mahdavi et al (2024) views inputs as context sequences. Kim et al (2023) and Kajitsuka and Sato (2024) prove memory capacity results in the sequence-to-sequence setting, where inputs are document sequences of token embeddings in $\mathbb{R}^{d\times T}$ and outputs are in $[C]^{1\times T}$ or $[C]^{d\times T}$ for some $C\in\mathbb{N}$. This does not apply to the next-token prediction setting for two reasons. First, the number of document sequences, $N$, could go to infinity while the number of unique contexts, $n$, remains fixed, thus the number of parameters required to apply the result, $\Omega(\sqrt{NT})$ in Kim et al (2023) and $\Omega(NT)$ in Kajitsuka and Sato (2024), is vacuous in the next-token prediction setting. Second, the outputs are not in $\mathbb{R}^{\omega\times T}$. Remark 3.2 of Kim et al (2023) explains how the codomain can be extended to ×$\mathbb{R}^{\ell\times T}$ for any $\ell\in\mathbb{N}$ by fixing the precision level, $\epsilon>0$, but the new required number of parameters is $\Omega(\ell\log(\ell/\epsilon)\sqrt{NT})$, which goes to infinity as $\epsilon$ goes to zero. Furthermore, both of these results assume that the embedded tokens are sufficiently separated while we do not make any assumptions on inputs.
Thus, the result of Mahdavi et al (2024) is the most relevant to ours. They prove a memory capacity result for transformers in the setting where inputs are context sequences of token embeddings in $\mathbb{R}^{d\times\tau}$ and each output is a prediction of the next-token embedding in $\mathbb{R}^{d}$, showing that $\approx nd$ parameters are sufficient to interpolate $n$ input-output pairs, which is optimal. Interestingly, they analyzed a model without a FNN sub-layer and instead leveraged the number of self-attention heads to prove their result. However, besides only applying to outputs in $\mathbb{R}^{d}$, rather than $\mathbb{R}^{\omega}$, their result requires that all context sequences have the same length.
In conclusion, while significant progress has been made in the analysis of the memory capacity of transformers in just the past year, our Theorems 5 is the first optimal memory capacity lower bound for transformers in the next-token prediction setting. Our upper bound, Theorem 1, can be seen as a generalization of Theorem 1 of Sontag (1997) from input-output pairs in $\mathbb{R}^{d}\times\mathbb{R}$ to input-output pairs in $\cup_{t=1}^{\infty}[\omega]^{t}\times\Delta^{\omega-1}$. However, our proof is quite different because $\cup_{t=1}^{\infty}[\omega]^{t}$ is countable.
As for Theorems 2 and 3, these are classical results when $a$ and $b$ are both generic, but to the best of our knowledge, they are novel results in the setting where $b\in\mathbb{R}^{n}$ has entries which are nonzero and distinct. However, closely related results have been shown before. Tamura (1991) proved a version of Theorem 3 when $\psi$ is the logistic function and restated the result in Tamura and Tateishi (1997). Then Huang and Babri (1998) proved it more generally when $\psi$ is bounded, nonlinear, and has a limit at either positive or negative infinity. Their result does not require $\psi$ to be real analytic at a point and not a polynomial there, but our result does not require $\psi$ to be bounded or to have limits. Thus, the two results are complementary. Interestingly, their proof sends the input values of $\psi$ off to infinity (or negative infinity), while our proof only uses input values in the interval of convergence of the point where $\psi$ is real analytic.
We have reviewed the works most closely related to our memory capacity results, but there are many relevant studies exploring the optimization of one-layer transformers as well, such as Tian et al (2023); Chen and Li (2024); Tarzanagh et al (2023); Zhang et al (2023). However, most do not address the next-token prediction scenario we study here. The exceptions, such as Makkuva et al (2024); Li et al (2024); Tian et al (2023, 2024); Ildiz et al (2024), rely on restrictive assumptions about input data and transformer dimensions, such as embedding size, to achieve minimal empirical risk. Additionally, optimization and learnability in the next-token prediction setting, albeit limited to linear models, have been recently studied in Malach (2023); Thrampoulidis (2024).
1.3 Organization
In Section 2 we present preliminaries necessary to prove our main results. In Section 3 we give a precise definition of memory capacity in the next-token prediction setting and prove the upper bound Theorem 1. In Section 4 we define the one-layer decoder-only transformer model. In Section 5 we prove the rank results. In Section 6 we prove the memory capacity results. In Section 7 we provide a numerical experiment demonstrating that not only is the memory capacity on the same order as the number of parameters, but the transformer can already be trained to the entropy lower bound in this regime. In Section 8, we conclude.
2 Preliminaries
Throughout the paper we use the following notation: $[n]$ denotes $\{1,\ldots,n\}$, $\Delta^{n-1}$ denotes the unit simplex in $\mathbb{R}^{n}$, $\mathrm{ri}\leavevmode\nobreak\ S$ denotes the relative interior of $S$, $\binom{A}{n}$ denotes $\{B\subset A\mid|B|=n\}$, vec denotes the column-wise vectorize operation, $e_{k}\in\mathbb{R}^{n}$ denotes the $k$th coordinate vector, $\mathbbm{1}_{n}$ denotes the vector of ones in $\mathbb{R}^{n}$, $I\in\mathbb{R}^{n\times n}$ denotes the identity matrix, $\delta$ denotes the Kronecker delta function, $a^{(k)}$ indicates that the exponent $k$ is applied to the vector $a$ element-wise, $\mathfrak{S}_{n}$ denotes the symmetric group of degree $n$, $\exp$ denotes the exponential function, and $\phi$ denotes the softmax function,
$\phi:\mathbb{R}^{n}\to\mathbb{R}^{n}:x\mapsto\frac{\exp(x)}{\sum_{i=1}^{n}\exp(x_{i})}.$
We also use the following conventions for matrices. If $a_{i}\in\mathbb{R}\leavevmode\nobreak\ \forall i\in[m]$, then $[a_{i}]_{i\in[m]}$ denotes $(a_{1},\ldots,a_{m})\in\mathbb{R}^{m}$. If $a_{j}\in\mathbb{R}^{m}\leavevmode\nobreak\ \forall j\in[n]$, then $[a_{j}]_{j\in[n]}$ denotes $[a_{1}|\cdots|a_{n}]\in\mathbb{R}^{m\times n}$. Going the other way, if $a\in\mathbb{R}^{m}$, then $a_{i}$ denotes element $i$ of $a$, and if $A\in\mathbb{R}^{m\times n}$, then $a_{i,j}$ denotes element $(i,j)$ of $A$. We will also index vectors and matrices using brackets. If $a\in\mathbb{R}^{m}$ and $\alpha\in[m]^{k}$, then $a[\alpha]$ denotes $[a_{\alpha_{i}}]_{i\in[k]}$ and $\mathrm{len}(a)$ denotes $m$. If $A\in\mathbb{R}^{m\times n}$, $\alpha\in[m]^{k}$, and $\beta\in[n]^{\ell}$, then $A[\alpha,\beta]$ denotes $[[a_{\alpha_{i},\beta_{j}}]_{i\in[k]}]_{j\in[\ell]}$. In this context, a colon denotes $(1,2,\ldots)$ and $i:j$ denotes $(i,\ldots,j)$.
In the realm of real analytic functions, we use Taylor’s theorem as as well as the fact that the zero set of a nontrivial real analytic function is measure zero and closed (Gunning and Rossi, 1965, Corollary 10). We say a property holds “generically” if the set of exceptions lies on the zero set of a nontrivial real analytic function. Note that $\phi$ is real analytic since the numerator and denominator are real analytic with the denominator being nowhere zero. Moreover, the entire self-attention sub-layer of the transformer model, which will be described in detail below, is real analytic since it is a composition of real analytic functions. Furthermore, if $\psi$ is real analytic at a point, then the entire transformer parameter mapping is real analytic at a point. Similarly, if $\psi$ is continuously differentiable, then the entire transformer parameter mapping is continuously differentiable, in which case we can apply Sard’s theorem (Sard, 1942).
We also use the model-theoretic structure $(\mathbb{R}_{\mathrm{an}},\exp)$ (van den Dries et al, 1994; van den Dries and Miller, 1994; van den Dries, 1998). Polynomials, restricted analytic functions, and exp are the basic $(\mathbb{R}_{\mathrm{an}},\exp)$-definable functions, but compositions, quotients (if the denominator is nowhere zero), inverses (of injective functions), and limits of $(\mathbb{R}_{\mathrm{an}},\exp)$-definable functions are $(\mathbb{R}_{\mathrm{an}},\exp)$-definable as well (van den Dries, 1998, Lma. 2.3). From this, it can be seen that $\phi$, tanh, arctan, erf, and GELU are $(\mathbb{R}_{\mathrm{an}},\exp)$-definable. Furthermore, if $\psi$ is $(\mathbb{R}_{\mathrm{an}},\exp)$-definable, then the entire transformer parameter mapping is $(\mathbb{R}_{\mathrm{an}},\exp)$-definable, and so its image either has nonempty interior or is nowhere dense (Sontag, 1996, Corollary B.4).
In the realm of linear algebra, we use the notion of Kruskal rank (Kruskal, 1977), as well as the Cauchy-Binet formula (Gantmacher, 1960, Sec. I.2.4) and the Leibniz determinant formula (Axler, 2015, Def. 10.33). In the realm of number theory, we use that $\exp(1)$ is transcendental (Hermite, 1873).
3 Next-token prediction
We precisely define the memory capacity of a model in the next-token prediction setting in the following way.
Definition 1.
Let $k,\omega,T\in\mathbb{N}$. For all $\theta\in\mathbb{R}^{k}$, let
$h_{\theta}:\cup_{t=1}^{T}[\omega]^{t}\to\mathbb{R}^{\omega}.$
Let $\phi$ be the softmax function. Then the memory capacity of $\{\phi\circ h_{\theta}\mid\theta\in\mathbb{R}^{k}\}$ is the largest $n\in\mathbb{N}$ such that for all $\tau_{1},\ldots,\tau_{n}\in[T]$ and distinct $\alpha^{1}\in[\omega]^{\tau_{1}},\ldots,\alpha^{n}\in[\omega]^{\tau_{n}}$,
$\{[(\phi\circ h_{\theta})(\alpha^{1})|\cdots|(\phi\circ h_{\theta})(\alpha^{n})]\mid\theta\in\mathbb{R}^{k}\}=(\mathrm{ri}\Delta^{\omega-1})^{n}.$
We call $[\omega]$ the “vocabulary” and its elements “tokens.” The tokens may represent letters, words, or something in between, as in Byte Pair Encoding (Gage, 1994). Given a sequence of tokens $\alpha\in[\omega]^{\tau}$, we think of $(\phi\circ h_{\theta})(\alpha)\in\mathrm{ri}\Delta^{\omega-1}$ as the model’s prediction of the conditional probability distribution of the token to follow $\alpha$. If $\theta\mapsto h_{\theta}(\alpha)$ is continuously differentiable for all $\alpha\in\cup_{t=1}^{T}[\omega]^{t}$, then we have the following upper bound on the memory capacity.
Lemma 1.
Let $k,\omega,T\in\mathbb{N}$. For all $\theta\in\mathbb{R}^{k}$, let
$h_{\theta}:\cup_{t=1}^{T}[\omega]^{t}\to\mathbb{R}^{\omega}.$
Assume that $\theta\mapsto h_{\theta}(\alpha)$ is continuously differentiable for all $\alpha\in\cup_{t=1}^{T}[\omega]^{t}$. Let $\phi$ be the softmax function. Then the memory capacity of $\{\phi\circ h_{\theta}\mid\theta\in\mathbb{R}^{k}\}$ is less than or equal to $k/(\omega-1)$.
Proof.
Let $\alpha^{1},\ldots,\alpha^{n}\in\cup_{t=1}^{T}[\omega]^{t}$ be distinct. Define $F$ as the mapping $\theta\mapsto[(\phi\circ h_{\theta})(\alpha^{1})|\cdots|(\phi\circ h_{\theta})(\alpha^{n})]$. $F$ is continuously differentiable, so we can apply Sard’s theorem (Sard, 1942). Note that $\mathrm{ri}\Delta^{\omega-1}$ is an $(\omega-1)$-dimensional manifold. Thus, if $k<n(\omega-1)$, then the image of $F$ has measure zero, hence $F$ is not surjective. ∎
In practice, we are generally given $N$ sequences $\beta^{1},\ldots,\beta^{N}\in[\omega]^{T+1}$, and $\theta$ is trained to minimize the cross-entropy loss,
$\mathcal{L}(\theta;\beta^{1},\ldots,\beta^{N})\coloneqq-\sum_{i=1}^{N}\sum_{t=1}^{T}\log\left(\phi\left(h_{\theta}\left(\beta_{1}^{i},\ldots,\beta_{t}^{i}\right)\right)_{\beta_{t+1}^{i}}\right).$
Enumerate the multiset $\{(\beta_{1}^{i},\ldots,\beta_{t}^{i})\mid i\in[N],t\in[T]\}$ as $\{(\alpha^{1},c_{1}),\ldots,(\alpha^{n},c_{n})\}$. Let $\tau_{i}$ denote the length of $\alpha^{i}$ for each $i\in[n]$ and define
$\hat{\pi}_{\gamma,i}\coloneqq\frac{|\{j\in[N]\mid(\beta_{1}^{j},\ldots,\beta_{\tau_{i}}^{j})=\alpha^{i},\beta_{\tau_{i}+1}^{j}=\gamma\}|}{c_{i}}\leavevmode\nobreak\ \forall(\gamma,i)\in[\omega]\times[n].$ (1)
For each $i\in[n]$, define $\hat{\pi}_{i}=[\hat{\pi}_{\gamma,i}]_{\gamma\in[\omega]}$. We call the $\alpha^{1},\ldots,\alpha^{n}$ the “contexts” of $\beta^{1},\ldots,\beta^{N}$. Then $c_{i}$ is the number of times that context $\alpha^{i}$ appears and $\hat{\pi}_{i}$ is the empirical distribution of the next-token conditioned on $\alpha^{i}$. Define the entropy
$\mathcal{H}(\beta^{1},\ldots,\beta^{N})\coloneqq-\sum_{i=1}^{n}\sum_{\gamma\leavevmode\nobreak\ :\leavevmode\nobreak\ \hat{\pi}_{\gamma,i}\neq 0}c_{i}\hat{\pi}_{\gamma,i}\log\left(\hat{\pi}_{\gamma,i}\right).$
Then the following lemma follows from Gibb’s inequality.
Lemma 2.
Let $k,\omega,T,N\in\mathbb{N}$. For all $\theta\in\mathbb{R}^{k}$, let
$h_{\theta}:\cup_{t=1}^{T}[\omega]^{t}\to\mathbb{R}^{\omega}.$
Let $\beta^{1},\ldots,\beta^{N}\in[\omega]^{T+1}$. Then,
$\mathcal{L}(\theta;\beta^{1},\ldots,\beta^{N})\geq\mathcal{H}(\beta^{1},\ldots,\beta^{N})\leavevmode\nobreak\ \forall\theta\in\mathbb{R}^{k}.$
Moreover, if $\{[(\phi\circ h_{\theta})(\alpha^{1})|\cdots|(\phi\circ h_{\theta})(\alpha^{n})]\mid\theta\in\mathbb{R}^{k}\}=(\mathrm{ri}\Delta^{\omega-1})^{n}$, then
$\inf_{\theta\in\mathbb{R}^{k}}\mathcal{L}(\theta;\beta^{1},\ldots,\beta^{N})=\mathcal{H}(\beta^{1},\ldots,\beta^{N}).$
Applying Lemma 2 to a memory capacity lower bound results in an “empirical memory capacity” lower bound, which can be compared to the following empirical memory capacity upper bound. Note that we have to further assume that $\theta\mapsto h_{\theta}(\alpha)$ is $(\mathbb{R}_{\mathrm{an}},\exp)$-definable for all $\alpha\in\cup_{t=1}^{T}[\omega]^{t}$. As an example, this is satisfied by the transformer model described in the next section for activations such as tanh, arctan, and GELU.
Theorem 1.
Let $k,\omega\in\mathbb{N}$. Set $n^{\prime}=\lfloor k/(\omega-1)\rfloor+1$ and $T=\lceil\log(n^{\prime})/\log(\omega)\rceil$. For all $\theta\in\mathbb{R}^{k}$, let
$h_{\theta}:\cup_{t=1}^{T}[\omega]^{t}\to\mathbb{R}^{\omega}.$
Assume that $\theta\mapsto h_{\theta}(\alpha)$ is continuously differentiable and $(\mathbb{R}_{\mathrm{an}},\exp)$-definable for all $\alpha\in\cup_{t=1}^{T}[\omega]^{t}$. Then there exists $N\geq n^{\prime}$ and $\beta^{1},\ldots,\beta^{N}\in[\omega]^{T+1}$ such that the number of contexts is less than or equal to $(2+1/(\omega-1))k/(\omega-1)+1$ and
$\inf_{\theta\in\mathbb{R}^{k}}\mathcal{L}(\theta;\beta^{1},\ldots,\beta^{N})>\mathcal{H}(\beta^{1},\ldots,\beta^{N}).$
Proof.
The size of $[\omega]^{T}$ is $\omega^{T}\geq n^{\prime}$, so there exists a distinct list $\alpha^{1},\ldots,\alpha^{n^{\prime}}\in[\omega]^{T}$. Choose such an $\alpha^{1},\ldots,\alpha^{n^{\prime}}$. Let $\beta^{1},\ldots,\beta^{n^{\prime}}\in[\omega]^{T+1}$ such that $\beta^{i}[1:T]=\alpha^{i}\leavevmode\nobreak\ \forall i\in[n^{\prime}]$. Then we can bound the total number of contexts
$n$ $\coloneqq n^{\prime}+\sum_{t=1}^{T-1}\omega^{t}$
$=n^{\prime}+\frac{\omega(\omega^{T-1}-1)}{\omega-1}$
$\leq n^{\prime}+\frac{\omega(n^{\prime}-1)}{\omega-1}$
$\leq\left(2+\frac{1}{\omega-1}\right)\frac{k}{\omega-1}+1.$
Define $F:\mathbb{R}^{k}\to(\mathrm{ri}\Delta^{\omega-1})^{n^{\prime}}:\theta\mapsto[(\phi\circ h_{\theta})(\alpha^{1})|\cdots|(\phi\circ h_{\theta})(\alpha^{n^{\prime}})]$ where $\phi$ is the softmax function. The softmax function is real analytic and $(\mathbb{R}_{\mathrm{an}},\exp)$-definable so $F$ is continuously differentiable and $(\mathbb{R}_{\mathrm{an}},\exp)$-definable. Since $k<n^{\prime}(\omega-1)$, the image of $F$ is measure zero by Sard’s theorem, hence has empty interior. So, by Corollary B.4 of Sontag (1996), it is nowhere dense, hence not dense. Let $Y\in(\Delta^{\omega-1})^{n^{\prime}}\backslash\overline{F(\mathbb{R}^{k})}$. Note that we can include more sequences $\beta^{n^{\prime}+1},\ldots,\beta^{N}$ without adding new contexts. Furthermore, we have control over the last token in every sequence, so by increasing $N$, we can make $[\hat{\pi}_{1}|\cdots|\hat{\pi}_{n^{\prime}}]$, as defined in Eq. (1), approximate $Y$ to arbitrary precision. Thus, since $Y\notin\overline{F(\mathbb{R}^{k})}$, there exists $N\geq n^{\prime}$ and $\beta^{1},\ldots,\beta^{N}\in[\omega]^{T+1}$ for which the conclusion holds. ∎
4 Transformer model
The one-layer transformer model is parameterized by the following matrices and vectors:
  • Let $Z\in\mathbb{R}^{d\times\omega}$. This is the “embedding matrix.” Its $\gamma$th column corresponds to the embedding of token $\gamma$ in $\mathbb{R}^{d}$. We call $d$ the “embedding dimension.”
  • Let $U\in\mathbb{R}^{d\times T}$. This is the “position matrix.” While we think of it as having only $T$ columns, it can actually be extended to arbitrarily many columns by appending new columns whenever an input with length greater than $T$ is encountered.
  • Let $W_{1,r},W_{2,r}\in\mathbb{R}^{d\times d_{r}},W_{3,r}\in\mathbb{R}^{d\times d_{0}}\leavevmode\nobreak\ \forall r\in[m_{0}],W_{0}\in\mathbb{R}^{m_{0}d_{0}\times d}$. These are the “self-attention matrices.” We call $m_{0}$ the “number of heads.”
  • Let $W\in\mathbb{R}^{d\times m},b\in\mathbb{R}^{m},V\in\mathbb{R}^{m\times\omega}$. These are the “hidden matrix,” “bias vector,” and “output matrix” respectively. We call $m$ the “number of neurons.”
Additionally, let $\psi:\mathbb{R}\to\mathbb{R}$ be real analytic at some point $\eta\in\mathbb{R}$ and not a polynomial there. This is the “activation function.” For example, $\psi$ could be a sigmoid, such as the logistic function, tanh, or arctan, or it could be a smoothed rectified linear unit, such as GELU.
Let $\theta$ denote the concatenation of all the model parameters. Then, for all $\tau\in\mathbb{N}$ and $\alpha\in[\omega]^{\tau}$, define
$\begin{split}&h_{1,\theta}(\alpha)=Z[:,\alpha]+U[:,1:\tau]\in\mathbb{R}^{d\times\tau}\\ &h_{2,\theta}(\alpha)=W_{0}^{\top}\text{vec}\left(\left[W_{3,r}^{\top}h_{1,\theta}(\alpha)\phi\left(h_{1,\theta}(\alpha)^{\top}W_{1,r}W_{2,r}^{\top}h_{1,\theta}(\alpha)[:,\tau]\right)\right]_{r\in[m_{0}]}\right)\in\mathbb{R}^{d}\\ &h_{\theta}(\alpha)=V^{\top}\psi\left(W^{\top}h_{2,\theta}(\alpha)+b\right)\in\mathbb{R}^{\omega}.\end{split}$ (2)
These three functions are the “embedding,” “self-attention,” and “FNN” sub-layers respectively of a single transformer layer. Note that we do not include the skip-connection $+h_{1,\theta}(\alpha)[\tau]$ in the self-attention sub-layer. This was omitted for simplicity, but its inclusion does not affect Theorems 4 and 5, which follow by almost the same proof.
5 The rank results
First, we need to derive the generic rank of $\psi(ab^{\top})$. We will do this for polynomial $\psi$ and then extend to $\psi$ which are real analytic at zero.
Theorem 2.
Let $K\subset\mathbb{N}\cup\{0\}$ and $c_{k}\in\mathbb{R}\backslash\{0\}\leavevmode\nobreak\ \forall k\in K$. Define $\psi(x)=\sum_{k\in K}c_{k}x^{k}$. Let $b\in\mathbb{R}^{n}$ have entries which are nonzero and distinct. Then the rank and Kruskal rank of $\psi(ab^{\top})$ are $\min\{m,n,|K|\}$ for generic $a\in\mathbb{R}^{m}$.
Proof.
Observe,
$\psi\left(ab^{\top}\right)$ $=\sum_{k\in K}c_{k}\left(ab^{\top}\right)^{(k)}=\sum_{k\in K}c_{k}a^{(k)}b^{(k)\top}.$
Let $s=\min\{m,n,|K|\}$. Let $I\subset[m]$ and $J\subset[n]$ such that $|I|=|J|=s$. Enumerate $I$ as $\{i_{1},\ldots,i_{s}\}$ and $J$ as $\{j_{1},\ldots,j_{s}\}$. We want to show that the $(I,J)$ minor is not identically zero. By the Cauchy-Binet formula,
$p(a)$ $\coloneqq\det_{I,J}\left(\psi\left(ab^{\top}\right)\right)=\sum_{S\in\binom{K}{s}}\left(\prod_{k\in S}c_{k}\right)\underset{\coloneqq p_{S}(a)}{\underbrace{\det_{I,[s]}\left(\left[a^{(k)}\right]_{k\in S}\right)}}\underset{\coloneqq q_{S}}{\underbrace{\det_{J,[s]}\left(\left[b^{(k)}\right]_{k\in S}\right)}}.$
Note that this would be identically zero if $s$ were greater than $|K|$ (since it would be an empty sum), thus the rank and Kruskal rank of $\psi(ab^{\top})$ are less than or equal to $\min\{m,n,|K|\}$ for all $a\in\mathbb{R}^{m}$. To go about proving the lower bound, first let $S\in\binom{K}{s}$ and enumerate it as $\{k_{1},\ldots,k_{s}\}$. Then, applying the Leibniz determinant formula, we get
$p_{S}(a)$ $=\sum_{\sigma\in\mathfrak{S}(s)}\mathrm{sgn}(\sigma)\underset{\coloneqq p_{S,\sigma}(a)}{\underbrace{a_{i_{1}}^{k_{\sigma(1)}}\cdots a_{i_{s}}^{k_{\sigma(s)}}}}.$
Let $\tau\in\mathfrak{S}(s)$. If $p_{S,\sigma}=p_{S,\tau}$, then $k_{\sigma(t)}=k_{\tau(t)}\leavevmode\nobreak\ \forall t\in[s]$, so $\sigma=\tau$, and so, collecting monomial terms, $p_{S,\sigma}$ has coefficient $\mathrm{sgn}(\sigma)$ in $p_{S}$. Let $S^{\prime}\in\binom{K}{s}$. If $p_{S,\sigma}=p_{S^{\prime},\tau}$, then, similarly, $S=S^{\prime}$ and $\sigma=\tau$. Thus, if $S^{\prime}\neq S$, then the monomial $p_{S,\sigma}$ has coefficient zero in $p_{S^{\prime}}$. So $(p_{S})$ is linearly independent. And the $q_{S}$ are nonzero since they are minors of a Vandermonde matrix with roots which are nonzero and distinct, so $p\not\equiv 0$, completing the proof. ∎
Theorem 3.
Let $\psi:\mathbb{R}\to\mathbb{R}$ be real analytic at zero and not a polynomial there. Let its radius of convergence at zero be $\rho$. Let $b\in\mathbb{R}^{n}$ have entries which are nonzero and distinct. Define $M=\{a\in\mathbb{R}^{m}\mid|a_{i}b_{j}|<\rho\leavevmode\nobreak\ \forall(i,j)\}$. Then $M$ is open and the rank and Kruskal rank of $\psi(ab^{\top})$ are $\min\{m,n\}$ for generic $a\in M$.
Proof.
To see that $M$ is open, notice that it is the pre-image of $(-\rho,\rho)^{m\times n}$ under the continuous map $a\mapsto ab^{\top}$. Now, let $(c_{k})$ be the coefficients of the Taylor expansion at zero of $\psi$. Define $K=\{k\in\mathbb{N}\cup\{0\}\mid c_{k}\neq 0\}$. Define $s=\min\{m,n\}$ and let $k_{1},\ldots,k_{s}$ be the first $s$ integers in $K$. Define $K_{s}=\{k_{1},\ldots,k_{s}\}$. Let $I\subset[m]$ and $J\subset[n]$ such that $|I|=|J|=s$. Define $p:M\to\mathbb{R}:a\mapsto\det_{I,J}(\sum_{k\in K_{s}}c_{k}(ab^{\top})^{(k)})$ and $q:M\to R:a\mapsto\det_{I,J}(\psi(ab^{\top}))$. Note from the proof of Theorem 2 that the monomials of $p$ all have degree $k_{1}+\cdots+k_{s}$ whereas if $S\in\binom{K}{s}$ has any integers not in $K_{s}$, then its corresponding monomials will have degree $>k_{1}+\cdots+k_{s}$. Thus, $p$ is the truncation of the Taylor expansion at zero of $q$ to precisely the monomials of degree $k_{1}+\cdots+k_{s}$. Since, by Theorem 2, $p$ is not identically zero, it has at least one nonzero coefficient, and so the Taylor expansion at zero of $q$ has at least one nonzero coefficient. Thus, $q$ is not identically zero, completing the proof. ∎
6 The memory capacity results
With Theorem 3 in hand, we are ready to prove the following fact about self-attention with embedding dimension 1.
Theorem 4.
Let $\omega,\tau\in\mathbb{N}$. Let $\phi$ be the softmax function. Define
$f:\leavevmode\nobreak$ $\mathbb{R}^{\omega}\times\mathbb{R}^{\tau}\times\cup_{t=1}^{\tau}[\omega]^{t}\to\mathbb{R}$
$(z,u,\alpha)\mapsto\langle x,\phi(xx_{t})\rangle$
$\text{where }t=\mathrm{len}(\alpha),x=z[\alpha]+u[1:t].$
Then $f(z,u,\cdot)$ maps injectively into \$\0\mathbb{R}\backslash\{0\}$ for generic $(z,u)$.
Proof.
First, let $A$ denote $\cup_{t=1}^{\tau}[\omega]^{t}$. Then, note that, for all $z\in\mathbb{R}^{\omega},u\in\mathbb{R}^{\tau},\sigma\in[\tau],\alpha\in[\omega]^{\sigma}$,
$f(z,u,\alpha)=\frac{\sum_{t=1}^{\sigma}(z_{\alpha_{t}}+u_{t})\exp((z_{\alpha_{t}}+u_{t})(z_{\alpha_{\sigma}}+u_{\sigma}))}{\sum_{t=1}^{\sigma}\exp((z_{\alpha_{t}}+u_{t})(z_{\alpha_{\sigma}}+u_{\sigma}))}.$
Clearly, $f(\cdot,\cdot,\alpha)$ is real analytic for all $\alpha\in A$, hence
$g\coloneqq\left(\prod_{\alpha\in A}f(\cdot,\cdot,\alpha)\right)\left(\prod_{\alpha,\beta\in A:\alpha\neq\beta}(f(\cdot,\cdot,\alpha)-f(\cdot,\cdot,\beta))\right)$
is real analytic. Note that the set of $(z,u)$ for which $f(z,u,\cdot)$ does not map injectively into \$\0\mathbb{R}\backslash\{0\}$ is precisely the zero set of $g$. Thus, the result follows if and only if $g$ is nontrivial if and only if (1) each $f(\cdot,\cdot,\alpha)$ is nontrivial and (2) each $f(\cdot,\cdot,\alpha)-f(\cdot,\cdot,\beta)$ is nontrivial. To prove (1), observe that $f(\mathbbm{1}_{\omega},0,\alpha)=1\leavevmode\nobreak\ \forall\alpha\in A$. Now, we will go about proving (2). First, let $\alpha,\beta\in A$. We want to show that $\alpha\neq\beta$ implies $f(\cdot,\cdot,\alpha)-f(\cdot,\cdot,\beta)$ is nontrivial. We will prove the contrapositive. Suppose $f(\cdot,\cdot,\alpha)-f(\cdot,\cdot,\beta)\equiv 0$. First, we will show that $\alpha$ and $\beta$ have the same length, then we will show that each entry of $\alpha-\beta$ is zero.
To show that $\alpha$ and $\beta$ have the same length, set $z=0$ and $u=(1,\ldots,\tau)$, and define the polynomials
$p_{\gamma}=\sum_{t=1}^{\mathrm{len}(\gamma)}tx^{t\mathrm{len}(\gamma)}\text{ and }q_{\gamma}=\sum_{t=1}^{\mathrm{len}(\gamma)}x^{t\mathrm{len}(\gamma)}\leavevmode\nobreak\ \forall\gamma\in\{\alpha,\beta\}.$
Note that the monomial $x^{\mathrm{len}(\alpha)^{2}+\mathrm{len}(\beta)^{2}}$ has coefficient $\mathrm{len}(\alpha)-\mathrm{len}(\beta)$ in $p_{\alpha}q_{\beta}-p_{\beta}q_{\alpha}$. Also, $f(z,u,\gamma)=p_{\gamma}(\exp(1))/q_{\gamma}(\exp(1))\leavevmode\nobreak\ \forall\gamma\in\{\alpha,\beta\}$. So, since $f(\cdot,\cdot,\alpha)-f(\cdot,\cdot,\beta)\equiv 0$, $(p_{\alpha}q_{\beta}-p_{\beta}q_{\alpha})(\exp(1))=0$. But, $\exp(1)$ is transcendental, so $p_{\alpha}q_{\beta}-p_{\beta}q_{\alpha}\equiv 0$. Thus, $\mathrm{len}(\alpha)=\mathrm{len}(\beta)$.
Let $\sigma=\mathrm{len}(\alpha)=\mathrm{len}(\beta)$. Let $s\in[\sigma]$. To show that $\alpha_{s}=\beta_{s}$, set $z=(1,\ldots,\omega)$ and $u=\omega e_{s}$, and define the polynomials
$p_{\gamma}=\sum_{t=1}^{\sigma}(\gamma_{t}+\omega\delta_{t,s})x^{(\gamma_{t}+\omega\delta_{t,s})(\gamma_{\sigma}+\omega\delta_{\sigma,s})}\text{ and }q_{\gamma}=\sum_{t=1}^{\sigma}x^{(\gamma_{t}+\omega\delta_{t,s})(\gamma_{\sigma}+\omega\delta_{\sigma,s})}\leavevmode\nobreak\ \forall\gamma\in\{\alpha,\beta\}.$
Note that the monomial $x^{(\alpha_{s}+\omega)(\alpha_{\sigma}+\omega\delta_{\sigma,s})+(\beta_{s}+\omega)(\beta_{\sigma}+\omega\delta_{\sigma,s})}$ has coefficient $\alpha_{s}-\beta_{s}$ in $p_{\alpha}q_{\beta}-p_{\beta}q_{\alpha}$. Also, $f(z,u,\gamma)=p_{\gamma}(\exp(1))/q_{\gamma}(\exp(1))\leavevmode\nobreak\ \forall\gamma\in\{\alpha,\beta\}$. So, again, $p_{\alpha}q_{\beta}-p_{\beta}q_{\alpha}\equiv 0$. Thus, $\alpha_{s}=\beta_{s}$. Since this was for arbitrary $s\in[\sigma]$, we get that $\alpha=\beta$, proving the theorem. ∎
It is natural to wonder what would happen if we replaced $\phi$ with the function $x\mapsto a^{x}/\sum_{\gamma=1}^{\omega}a^{x_{\gamma}}$ for some $a>1$ which is not transcendental. In this case, since $a^{b}=\exp(b\log(a))$, we would just need to scale our example $(z,u)$’s by $\sqrt{\log(a)}$ to get the same memory capacity result.
Combining Theorem 3 and Theorem 4 leads to a memory capacity lower bound for one-layer decoder-only transformers.
Theorem 5.
Let $d,m_{0},d_{0},m,\omega,T\in\mathbb{N}$ and $d_{r}\in\mathbb{N}\leavevmode\nobreak\ \forall r\in[m_{0}]$. Let $\phi$ be the softmax function and let $\psi:\mathbb{R}\to\mathbb{R}$ be real analytic at some $\eta\in\mathbb{R}$ and not a polynomial there. Let $h_{\theta}$ be the one-layer transformer model defined in Equation (LABEL:eq:model). Then the memory capacity of $\{\phi\circ h_{\theta}\mid\theta\}$ is greater than or equal to $m$.
Proof.
We will reduce to the setting where $d=m_{0}=d_{0}=d_{1}=1$. Given $z\in\mathbb{R}^{\omega},u\in\mathbb{R}^{T},w\in\mathbb{R}^{m}$, define $Z=\mathbbm{1}_{d}z^{\top}/\sqrt{d},U=\mathbbm{1}_{d}u^{\top}/\sqrt{d},W_{1,r}=W_{2,r}=\mathbbm{1}_{d}\mathbbm{1}_{d_{r}}^{\top}/\sqrt{dd_{r}}\leavevmode\nobreak\ \forall r\in[m_{0}],W_{3,r}=\mathbbm{1}_{d}\mathbbm{1}_{d_{0}}^{\top}/\sqrt{dd_{0}}\leavevmode\nobreak\ \forall r\in[m_{0}],W_{0}=\mathbbm{1}_{m_{0}d_{0}}\mathbbm{1}_{d}^{\top}/(m_{0}\sqrt{dd_{0}}),W=\mathbbm{1}_{d}w^{\top}/\sqrt{d},b=\eta\mathbbm{1}_{m}$. Then, for all $\alpha\in\cup_{t=1}^{T}[\omega]^{t}$, $W^{\top}h_{2,\theta}(\alpha)=wf(z,u,\alpha)$, where we use the definition of $f$ from Theorem 4.
Set $n=m$. Let $\alpha^{1},\ldots,\alpha^{n}\in\cup_{t=1}^{T}[\omega]^{t}$ be distinct. Invoking Theorem 4, choose $(z,u)$ such that $[f(z,u,\alpha^{i})]_{i\in[n]}$ has entries which are nonzero and distinct. Invoking Theorem 3, choose $w$ such that $\psi(w[f(z,u,\alpha^{i})]_{i\in[n]}^{\top}+b\mathbbm{1}_{n}^{\top})$ has rank $n$. Let $Y\in\mathbb{R}^{\omega\times n}$. Then, there exists $V\in\mathbb{R}^{m\times\omega}$ such that $Y=V^{\top}\psi(w[f(z,u,\alpha^{i})]_{i\in[n]}^{\top}+b\mathbbm{1}_{n}^{\top})$. Since $\phi(\mathbb{R}^{\omega})=\mathrm{ri}\Delta^{\omega-1}$, the result follows. ∎
In the proof, token $\gamma\in[\omega]$ is embedded as $z_{\gamma}\mathbbm{1}_{d}/\sqrt{d}$ and the $t$th position vector is $u_{t}\mathbbm{1}_{d}/\sqrt{d}$. Thus, the inputs to the self-attention sub-layer are sequences of vectors coming from a one-dimensional subspace, so it makes sense for the self-attention matrices to be constant matrices. Finally, the $k$th hidden neuron of the FNN sub-layer has weight vector $w_{k}\mathbbm{1}_{d}/\sqrt{d}$.
Theorem 5 also suggests an algorithm for finding an interpolating solution. First, note that every minor of $\psi(W^{\top}[h_{2,\theta}(\alpha^{i})]_{i\in[n]}+b\mathbbm{1}_{n}^{\top})$ is real analytic on an open neighborhood of zero. Thus, it works to (1) sample $\theta$ from any continuous probability distribution, (2) choose $\epsilon>0$ sufficiently small so that the entries of $\epsilon W^{\top}[h_{2,\theta}(\alpha^{i})]_{i\in[n]}+b\mathbbm{1}_{n}^{\top}$ are each in the interval of convergence of $\psi$, and (3) solve for $V$ in $Y=V^{\top}\psi(W^{\top}[h_{2,\theta}(\alpha^{i})]_{i\in[n]}+b\mathbbm{1}_{n}^{\top})$.
To compare the lower bound of Theorem 5 to the upper bound of Lemma 1, let $d_{r}=d_{1}\leavevmode\nobreak\ \forall r\in[m_{0}]$ for simplicity. Then, the total number of parameters is $\omega m+m(d+1)+2m_{0}(d_{0}+d_{1})d+(\omega+T)d$. Thus, the ratio between the upper and lower bounds is
$1+\frac{d+2}{\omega-1}+\frac{2m_{0}(d_{0}+d_{1})d}{(\omega-1)m}+\frac{(\omega+T)d}{(\omega-1)m},$ (3)
which is $O(1)$ if and only if the number of parameters in the FNN is at least a constant proportion of the total number of parameters. To see that this is generally the case, consider how, in practice, $\min\{\omega,m\}\geq d\geq\max\{d_{0},d_{1}\}$ and $\omega m\geq Td$, so the ratio reduces to $O(m_{0}d^{2}/(\omega m))$, which is more clearly $O(1)$.
Because the only two properties of self-attention that were used in the proof of Theorem 5 were that (1) $f(z,u,\cdot)$ maps injectively into \$\0\mathbb{R}\backslash\{0\}$ for generic $(z,u)$ and (2) we can reduce to the $d=1$ case, it is natural to wonder what alternatives to self-attention share this property. As it turns out, it works to simply take the weighted average of embedded vectors, a mechanism that we call “token averaging.”
Theorem 6.
Let $\omega,\tau\in\mathbb{N}$. Define
$f:\leavevmode\nobreak$ $\mathbb{R}^{\omega}\times\mathbb{R}^{\tau}\times\cup_{t=1}^{\tau}[\omega]^{t}\to\mathbb{R}$
$(z,u,\alpha)\mapsto\sum_{t=1}^{\mathrm{len}(\alpha)}u_{t}z_{\alpha_{t}}.$
Then $f(z,u,\cdot)$ maps injectively into \$\0\mathbb{R}\backslash\{0\}$ for generic $(z,u)$.
Proof.
First, let $A$ denote $\cup_{t=1}^{\tau}[\omega]^{t}$ and define
$g=\left(\prod_{\alpha\in A}f(\cdot,\cdot,\alpha)\right)\left(\prod_{\alpha,\beta\in A:\alpha\neq\beta}(f(\cdot,\cdot,\alpha)-f(\cdot,\cdot,\beta))\right).$
Note that the set of $(z,u)$ for which $f(z,u,\cdot)$ does not map injectively into \$\0\mathbb{R}\backslash\{0\}$ is precisely the zero set of $g$. Thus, the result follows if and only if $g$ is nontrivial if and only if (1) each $f(\cdot,\cdot,\alpha)$ is nontrivial and (2) each $f(\cdot,\cdot,\alpha)-f(\cdot,\cdot,\beta)$ is nontrivial. Note that $f(\mathbbm{1}_{\omega},\mathbbm{1}_{\tau},\alpha)=\mathrm{len}(\alpha)$. This proves (1). It also proves (2) in the case where $\mathrm{len}(\alpha)\neq\mathrm{len}(\beta)$. All that is left is to prove (2) in the case where $\mathrm{len}(\alpha)=\mathrm{len}(\beta)$. Let $\alpha,\beta\in A$ such that $\alpha\neq\beta$ and $\sigma\coloneqq\mathrm{len}(\alpha)=\mathrm{len}(\beta)$. Then there exists $s\in[\sigma]$ such that $\alpha_{s}\neq\beta_{s}$. Choose such an $s$. Set $z=(1,\ldots,\omega)$ and $u=e_{s}$. Then $f(z,u,\alpha)=\alpha_{s}\neq\beta_{s}=f(z,u,\beta)$, proving (2), and so proving the theorem. ∎
Theorem 7.
Let $d,m,\omega,T\in\mathbb{N}$. For all $\theta\coloneqq(Z,u,W,b,V)\in\mathbb{R}^{d\times\omega}\times\mathbb{R}^{T}\times\mathbb{R}^{d\times m}\times\mathbb{R}^{m}\times\mathbb{R}^{m\times\omega}$, define
$h_{\theta}:\cup_{t=1}^{T}[\omega]^{t}$ $\to\mathbb{R}^{\omega}$
$\alpha$ $\mapsto V^{\top}\psi\left(W^{\top}\sum_{t=1}^{\mathrm{len}(\alpha)}u_{t}z_{\alpha_{t}}+b\right)$
where $\psi:\mathbb{R}\to\mathbb{R}$ is real analytic at some $\eta\in\mathbb{R}$ and not a polynomial there. Then the memory capacity of $\{\phi\circ h_{\theta}\mid\theta\}$ is greater than or equal to $m$.
Proof.
Set $n=m$. Let $\alpha^{1},\ldots,\alpha^{n}\in\cup_{t=1}^{T}[\omega]^{t}$. Invoking Theorem 6, choose $(z,u)$ such that $f(z,u,\cdot)$ maps injectively into \$\0\mathbb{R}\backslash\{0\}$. Then $[f(z,u,\alpha^{i})]_{i\in[n]}$ has entries which are nonzero and distinct. Set $b=\eta\mathbbm{1}_{m}$. Invoking Theorem 3, choose $w$ such that $\psi(w[f(z,u,\alpha^{i})]_{i\in[n]}^{\top}+b\mathbbm{1}_{n}^{\top})$ has rank $n$. Let $Y\in\mathbb{R}^{\omega\times n}$. Then there exists $V\in\mathbb{R}^{m\times\omega}$ such that $Y=V^{\top}\psi(w[f(z,u,\alpha^{i})]_{i\in[n]}^{\top}+b\mathbbm{1}_{n}^{\top})$. Choose such a $V$. Define $Z=\mathbbm{1}_{d}z^{\top}/\sqrt{d}$ and $W=\mathbbm{1}_{d}w^{\top}/\sqrt{d}$. Then $W^{\top}Z=wz^{\top}$, so $Y=[h_{\theta}(\alpha^{i})]_{i\in[n]}$. Finally, since $\phi(\mathbb{R}^{\omega})=\mathrm{ri}\Delta^{\omega-1}$, the result follows. ∎
7 Experiment
We also provide an experiment showing that $\Omega(n)$ parameters seems to be sufficient not only in terms of memory capacity, but also in terms of training. In Figure 1, it can be seen that as the width of the hidden dimension (referred to as $m$ in the paper) of our decoder-only transformer model increases, the training error decreases. Furthermore, the number of parameters required to reach the entropy lower bound seems to be roughly linear, suggesting that first-order optimization methods can find an interpolating solution almost as soon as one exists.
Figure 1: We show that our model requires more parameters to memorize an increasing number of unique contexts. Left: As the hidden dimension, $m$, increases and the number of unique contexts, $n$, decreases, the gap between the training error and the entropy lower bound trends downwards. Right: As the number of unique contexts increases, the minimum number of parameters required for the gap between the training error and the entropy lower-bound to fall below the minimum threshold increases.
All experiments are performed on the TinyStories data set (Eldan and Li, 2023), where we take subsets of size 100, 200, 300, 400, 500, 600, 700, 800, 900, 1000 from the data set. These data subsets have 377, 699, 952, 1123, 1459, 1637, 1896, 2117, 2374, 2637 unique contexts, respectively. Note that in our line plot, the minimum threshold varies for each data subset and is computed by taking the entropy lower bound and multiplying it by 0.01 (this helps to account for the fact that larger data sets have larger entropy lower bounds). We train 8 transformer models for each data subset, with $m$ values 4, 8, 16, 32, 64, 128, 256, 512. For simplicity and reduced computation time, we truncate each sequence to have a length of 10. Finally, we use an embedding size of 16 and the sinusoidal positional embedding from Vaswani et al (2017).
The implementation of our model is a decoder-only transformer and does not have any encoder blocks. It has only one decoder block, which consists of a self-attention layer followed by a linear layer, after which a GELU activation function is applied. Only one head is used for the self-attention. The output of this decoder block is fed into another linear layer, which produces the final output. We include a bias in both linear layers. Unlike many of the more complicated transformer models, we do not use any dropout, layer normalization, or residual connections in our implementation.
To compute the training error on each iteration, the cross-entropy loss function is used. For optimization, we use the Adam optimizer (Kingma and Ba, 2015) in full batch mode, with a stepsize of 0.0001 and no regularization (sometimes referred to as weight decay). Our training error values are computed based on 50,000 iterations. We choose a large number of iterations since our data subsets are small and we want to see how close we can get to the entropy lower bound for each value of $m$. Finally, we run our experiments on a computer with a Intel(R) Core(TM) i9-10900X CPU @ 3.70GHz processer and a NVIDIA RTX A4000 GPU with 16GB of memory. We refer you to the following Github link for the code to run our experiments: https://github.com/curtfox/decoder-memory-capacity.
8 Conclusion
We proved upper and lower bounds on the memory capacity of one-layer decoder-only transformers in the setting of next-token prediction. Assuming that the activation function is continuously differentiable and $(\mathbb{R}_{\mathrm{an}},\exp)$-definable (e.g. arctan, tanh, GELU), we showed that $\Omega(n\omega)$ parameters are necessary to memorize a data set with $n$ unique contexts and a vocabulary of size $\omega$. Assuming that the activation function is real analytic at a point and not a polynomial there (e.g. arctan, tanh, GELU), we showed that $\Omega(n\omega)$ parameters is also sufficient. Furthermore, we provided numerical evidence that the transformer can already be trained to the entropy lower bound in the $\Theta(n\omega)$ parameter regime. Theoretically investigating optimization and generalization in the $\Theta(n\omega)$ parameter regime is left as a future direction.